OpenStudy (anonymous):

consider the curve given by xy^2-x^3y=6 Show that dy/dx=3x^2y-y^2/2xy-x^3

OpenStudy (anonymous):

Use implicit derivation. Do you know how?

OpenStudy (anonymous):

No, sadly.

OpenStudy (anonymous):

You have to suppose that y = y(x) and you don't know, explicitly, its derivative. Hence, you differentiate as you always did, thus: xy^2-x^3y=6 <=> y^2 + x*2y*(dy/dx) + [rest] Is it clear?

OpenStudy (anonymous):

I don't understand the second part? y^2 + x*2y*(dy/dx) + [rest]

OpenStudy (anonymous):

If you had, instead, this function to differentiate this function : x^(3n) How would you do?

OpenStudy (anonymous):

I honestly have no idea. I missed this lesson during class.

OpenStudy (anonymous):

How can you miss the basics of derivative of a function like x^n and now you are asked to do implicit differentiation... I'm not a fool, sorry.

OpenStudy (anonymous):

x^n= nx^n-1